# SignX: Continuous Sign Recognition in Compact Pose-Rich Latent Space

### v2

@@ Abstract
The complexity of sign language data processing brings many challenges. The current approach to recognition of ASL signs aims to translate RGB sign language videos through pose information into English-based ID Glosses, which serve to uniquely identify ASL signs [^1] Note that there is no shared convention for assigning such glosses to ASL signs, so consistent glossing conventions must be used across all datasets.. This paper proposes SignX, a novel framework for continuous sign language recognition in compact pose-rich latent space. First, we construct a unified latent representation that encodes heterogeneous pose formats (SMPLer-X, DWPose, Mediapipe, PrimeDepth, and Sapiens Segmentation) into a compact, information-dense space. Second, we train a ViT-based Video2Pose module to extract this latent representation directly from raw videos. Finally, we develop a temporal modeling and sequence refinement method that operates entirely in this latent space. This multi-stage design achieves end-to-end sign language recognition while significantly reducing computational consumption. Experimental results demonstrate that SignX achieves state-of-the-art accuracy on continuous sign language recognition. More information at anonymous url.
@@ 1 Introduction
Figure 1: Multimodal pose estimation methods: SMPLer-X cai2023smplerx can provide accurate 3D human body parameters; DWPose yang2023effective focuses on real-time 2D keypoint detection; Mediapipe MediaPipe provides lightweight but efficient 3D pose prediction; PrimeDepth zavadski2024primedepth can obtain scene depth information; while Sapiens Segmentation khirodkar2024sapiens provides fine-grained human body part segmentation results. These methods each have their own characteristics, providing rich feature representations for sign language recognition.
@@
Sign Language Recognition (SLR) aims to automatically convert sign language videos into text or glosses (represented as upper-case text, which serve as unique identifiers of signs). It has important social value in facilitating barrier-free communication between deaf and hearing individuals camgoz2018neural; camgoz2020sign; stoll2020text2sign; yin-etal-2021-including-signed-languages. However, SLR faces technical challenges: (1) Signed languages are complex multimodal movements, involving hand and arm movements as well as facial expressions and body postures, with complex temporal dependencies among these modalities Bohacek_2022_WACV; signor; SIGNUM. (2) Existing SL datasets are limited in size and lack a uniform preprocessing format, with complicated processing pipelines, which severely constrains the application of deep learning methods duarte2021how2sign; JAsigning; camgoz2020sign. (3) Different pose processing formats attend to different visual aspects, leading to distributional shifts that limit effective integration of innovative advantages across model types fang2025signllmsignlanguageproduction; fang2025signdiffdiffusionmodelamerican.
@@
To address these challenges, we develop a Vid2Pose module by using a novelty ViT model dosovitskiy2020vit, which inspired by the five most powerful pose estimation models. As shown in Fig. 1, we learn the Vid2Pose process in the encoder of the ViT model through a unified pose representation. Specifically, SMPLer-X cai2023smplerx provides 3D body mesh with kinematic and pose parameters for modeling body dynamics; DWPose yang2023effective captures 2D keypoints with rich facial landmarks essential for grammatical expressions; MediaPipe MediaPipe offers 3D joint coordinates that represent overall body motion trends. Additionally, PrimeDepth zavadski2024primedepth provides spatial depth information for disambiguating front-back positioning, while Sapiens Segmentation khirodkar2024sapiens captures human body shape through fine-grained part boundaries. This complementary design captures not only hand movements but also fine-grained details such as body dynamics, facial expressions, motion trajectories, spatial depth, and body shape, with the goal of fully capturing a person’s pose saunders2020progressive; stoll2020text2sign; pose_format_helper. This end-to-end processing approach greatly simplifies the pose representation Bohacek_2022_WACV; cho2021unifying; chiang2019unified process that is necessary for SL recognition.
@@
To address these limitations, we develop a recognition framework that operates entirely in the compact pose-rich latent space. Starting from the 2048-dimensional pose features extracted by Vid2Pose, we employ a ResNet34 backbone 7780459 followed by TemporalConv layers 10.1007/978-3-319-49409-8_7 to capture hierarchical temporal patterns. These are jointly learned with the pose features. To refine these predictions into coherent sequences, we apply a Transformer-based encoder-decoder zhang2023sltunet that performs sequence-to-sequence refinement with beam search and CTC regularization camgoz2020sign; vaswani2017attention. This design achieves accurate recognition directly in latent space, eliminating the need for raw pose processing and reducing computational overhead compared to vision-based approaches.
Through this muti-stage training, SignX achieves powerful functionality while maintaining architectural simplicity: It learns from various formats of prior knowledge and can directly use the raw video as input. Experimental results show that SignX achieves good performance on our ASL and mainstream datasets asllrp2025signbank; forster2012rwth; Zhou2021ImprovingSLT-with-monolingual-CSLDaily; WLASL, surpassing existing methods in both accuracy and robustness Chen-arxiv-2023-Robust.
Figure 2: Overview of Constructing Compact Pose-Rich Latent Space: Overall, we utilize ViT dosovitskiy2020vit to construct and accommodate a pose latent space. It has two entry points: a video entry at the top layer and a pose data entry in the middle section. (a) For training stage 1, we first train the pose fusion layer to output simple text information, ensuring that the learned pose representations are meaningful. (b) For training stage 2, we freeze all other components and only learn how RGB videos can be correctly converted into our pose features. For inference, only RGB is used as input, so we must ensure that we can encode RGB inputs into our pose features.
@@ 2 Related Work
While traditional SLR frameworks predominantly rely on feature extraction from high-dimensional pixel streams or raw skeletal coordinates, these approaches are often hindered by data noise and the difficulty of capturing fine-grained grammatical nuances. In contrast, diffusion models have recently redefined sign language generation by operating within compressed, structured latent spaces fang2025stablesignerhierarchicalsign; fang2025signllmsignlanguageproduction. Departing from existing SLR paradigms, our work introduces a novel recognition framework that performs inference directly within a sign-specific latent space. Inspired by the high-efficiency latent modeling in diffusion processes rombach2021highresolution; fang2025streamflowtheoryalgorithmimplementation, we leverage a pose-rich latent domain to bypass the limitations of raw data processing. This shift allows our model to more effectively capture complex spatiotemporal dependencies and linguistic structures, establishing a new direction for robust and efficient American Sign Language recognition.
@@ 3.1 Data Processing
To address the challenge of integrating heterogeneous pose information for SL recognition, we first construct a comprehensive processing pipeline that standardizes multiple pose formats. First, we use a high-quality ASLLRP SignStream® 3 Corpus neidle2022asl; asllrp2025signbank dataset, which contains over 80 hours of American Sign Language (ASL) videos with synchronized front view, side view, and facial close-up recordings. For each input video $V\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the number of frames and $H,W$ represent height and width, respectively, we extract sequential pose representations for each pose type: $P_{\text{DWPose}}\in\mathbb{R}^{T\times 384}$, $P_{\text{MediaPipe}}\in\mathbb{R}^{T\times 258}$, $_{\text{SMPLer-X}}\in\mathbb{R}^{T\times 165}, PP$$_{\text{PrimeDepth}}\in\mathbb{R}^{T\times 576}$, $P_{\text{Sapiens}}\in\mathbb{R}^{T\times 576}$. The dimensionality of each pose type reflects its specific information structure. For instance, DWPose contains 18 body keypoints, 21 keypoints for each hand, and 68 facial keypoints, along with their respective confidence scores, totaling 384 dimensions per frame.
@@ 3.2 Step I: ViT for Multimodal Pose Fusion
We initially attempted to use T5 2020t5, currently one of the most popular text decoders, as our text decoding component. However, experimental results revealed that for sign language translation tasks, we only need to consider the vocabulary present in our specific dataset. Large-scale decoders like T5 not only reduce inference speed, but also increase error probability, potentially generating words that do not exist in our target vocabulary $\mathcal{V}_{\text{dataset}}$. Therefore, we developed a custom CodeBook using training data, where the model outputs vocabulary indices rather than raw text tokens, significantly improving accuracy.
Figure 3: Organize and conduct continuous recognition in the latent space: (a) For the latent space of the enriched poses, we further distill, compress and organize it. The number of features trained should be as aligned as possible with the number of Gloss. This can further enhance the upper limit and performance of our model. (b) Then, we develop it based on a BiLSTM method zhang2023sltunet, enabling it to perform continuous Sign recognition in this latent space, thereby achieving results far superior to previous works.
@@ 3.3 Step II: ViT for Video2Pose
To process raw video inputs directly, we develop a Video2Pose module based on video entry of ViT dosovitskiy2020vit. Given an input video $V\in\mathbb{R}^{T\times H\times W\times 3}$, the Video2Pose module extracts frame-level features and models their temporal relationships:
@@
where $\hat{P}_{i}$ is the predicted pose of type $i$. This end-to-end approach eliminates the need for separate pose extraction models during deployment, greatly simplifying the application process saunders2021mixed; slt-how2sign-wicv2023; chen2020simple; xie2021segformer; xie2021simmim.
@@ 4.1 Latent Feature Organization and Optimization
Before the temporal learner can reason over sequences, we enforce a disciplined organization of the features produced by the Vid2Pose encoder. Every frame-level embedding $z_{t}\in\mathbb{R}^{2048}$ is accompanied by its gloss supervision signal $g_{t}$ and confidence mask $m_{t}$, giving us a richer tuple $\langle z_{t},g_{t},m_{t}\rangle$. As shown in Fig. 3, this tuple is passed through a lightweight transformer stack that performs (i) variance-preserving layer normalization, (ii) cross-frame covariance whitening (using coefficient $\gamma$), and (iii) stochastic frame dropping (with probability $\rho$) to avoid overfitting to repetitive segments. The process aligns pose cues with textual semantics via a GlossAlign function this process is formalized in Algorithm 1. Here, GlossAlign assigns each feature to its corresponding gloss based on temporal overlap, and CrossCov zhang2023sltunet computes the cross-covariance between consecutive frames for redundancy detection.
@@
To combat data scarcity in sign language datasets, we apply a $N=10$-fold temporal augmentation during feature extraction zhang2023sltunet. For each video sample $i$, the SMKD model zhang2023sltunet generates $N$ augmented versions through (i) random temporal scaling $\sigma\in[0.8,1.2]$, (ii) spatial jittering with probability $p_{jit}=0.3$, and (iii) feature-space Gaussian noise $\mathcal{N}(0,0.01)$. The augmented features are stored in HDF5 format with keys $\{i\}_{j}$ where $j\in[0,N-1]$ indexes the augmentation version. It enables efficient random access during training while expanding the effective data size from $M$ to $M\times N$ samples. Unlike pixel-space augmentation which requires re-encoding, our latent augmentation operates on extracted features, reducing computational overhead.
@@ 4.2 Latent-Space Temporal Modeling and Decoding
Sign2Gloss C2SLR zuo2022c2slr 52.50 50.10 38.20 29.50 24.76 52.30 49.95 38.00 29.35 24.58
CorrNet+ hu2024corrnetsignlanguagerecognition 53.80 51.40 39.50 31.20 26.14 53.60 51.25 39.30 31.05 25.96
SLTUNET zhang2023sltunet 54.20 52.10 40.10 31.80 26.48 54.05 51.90 39.95 31.65 26.31
@@
Table 1: Benchmark on the ASLLRP dataset neidle2022asl; asllrp2025signbank. To ensure a fair and rigorous comparison on the ASLLRP dataset, we re-implemented and fine-tuned several representative SLR frameworks, including C2SLR and SLTUNET, achieving their best possible performance as modern baselines.
@@ 5 Experimental Results
Here, we evaluate our SignX continuous recognition method on mainstream datasets (ASLLRP asllrp2025signbank: 2,669 utterances; PHOENIX2014-T camgoz2018neural: 7,096/519/642 train/val/test; CSL-Daily zhou2021improving: 18,401/1,077/1,176; WLASL-2000 WLASL: 2,000 signs), using metrics such as Bleu bleu, WER KLAKOW200219, and P-I li2020word. We conduct a series of experiments including qualitative, quantitative, ablation, and efficiency evaluations, demonstrating that our method has achieved a significant improvement over the state-of-the-art [^2] More mertric information in Section D.1..
@@ 5.1 Qualitative Assessment
SLTUNET zhang2023sltunet 52.23 - - - 27.87 52.11 52.92 41.76 33.99 28.47
TwoStream-SLT chen2022two 54.08 54.32 41.99 34.15 28.66 53.48 54.90 42.43 34.46 28.95
CorrNet+ hu2024corrnetsignlanguagerecognition 54.54 54.56 42.31 34.48 29.13 53.76 55.32 42.74 34.86 29.42
@@
SLTUNET zhang2023sltunet 53.58 - - - 23.99 54.08 54.98 41.44 31.84 25.01
TwoStream-SLT chen2022two 55.10 55.21 42.31 32.71 25.76 55.72 55.44 42.59 32.87 25.79
CorrNet+ hu2024corrnetsignlanguagerecognition 55.52 55.64 42.78 33.13 26.14 55.84 55.82 42.96 33.26 26.14
Uni-Sign li2025uni 56.03 55.30 - - 26.25 56.51 55.08 - - 26.36

### v3

@@ Abstract
The complexity of sign language data processing brings many challenges. The current approach to recognition of ASL signs aims to translate RGB sign language videos through pose information into English-based ID Glosses, which serve to uniquely identify ASL signs [^1] Note that there is no shared convention for assigning such glosses to ASL signs, so consistent glossing conventions must be used across all datasets.. This paper proposes SignX, a novel framework for continuous sign language recognition in compact pose-rich latent space. First, we construct a unified latent representation that encodes heterogeneous pose formats (SMPLer-X, DWPose, Mediapipe, PrimeDepth, and Sapiens Segmentation) into a compact, information-dense space. Second, we train a ViT-based Video2Pose module to extract this latent representation directly from raw videos. Finally, we develop a temporal modeling and sequence refinement method that operates entirely in this latent space. This multi-stage design achieves end-to-end sign language recognition while significantly reducing computational consumption. Experimental results demonstrate that SignX achieves state-of-the-art accuracy on continuous sign language recognition.
@@ 1 Introduction
Figure 1: Multimodal pose estimation methods: SMPLer-X Cai et al. (2023) can provide accurate 3D human body parameters; DWPose Yang et al. (2023) focuses on real-time 2D keypoint detection; Mediapipe MediaPipe provides lightweight but efficient 3D pose prediction; PrimeDepth Zavadski et al. (2024) can obtain scene depth information; while Sapiens Segmentation Khirodkar et al. (2024) provides fine-grained human body part segmentation results. These methods each have their own characteristics, providing rich feature representations for sign language recognition.
@@
Sign Language Recognition (SLR) aims to automatically convert sign language videos into text or glosses (represented as upper-case text, which serve as unique identifiers of signs). It has important social value in facilitating barrier-free communication between deaf and hearing individuals camgoz2018neural; camgoz2020sign; stoll2020text2sign; yin-etal-2021-including-signed-languages. However, SLR faces technical challenges: (1) Signed languages are complex multimodal movements, involving hand and arm movements as well as facial expressions and body postures, with complex temporal dependencies among these modalities Boháček and Hrúz (2022); signor; SIGNUM. (2) Existing SL datasets are limited in size and lack a uniform preprocessing format, with complicated processing pipelines, which severely constrains the application of deep learning methods duarte2021how2sign; JAsigning; camgoz2020sign. (3) Different pose processing formats attend to different visual aspects, leading to distributional shifts that limit effective integration of innovative advantages across model types Fang et al. (2025a, c).
@@
To address these challenges, we develop a Vid2Pose module by using a novelty ViT model Dosovitskiy et al. (2021), which inspired by the five most powerful pose estimation models. As shown in Fig. 1, we learn the Vid2Pose process in the encoder of the ViT model through a unified pose representation. Specifically, SMPLer-X Cai et al. (2023) provides 3D body mesh with kinematic and pose parameters for modeling body dynamics; DWPose Yang et al. (2023) captures 2D keypoints with rich facial landmarks essential for grammatical expressions; MediaPipe MediaPipe offers 3D joint coordinates that represent overall body motion trends. Additionally, PrimeDepth Zavadski et al. (2024) provides spatial depth information for disambiguating front-back positioning, while Sapiens Segmentation Khirodkar et al. (2024) captures human body shape through fine-grained part boundaries. This complementary design captures not only hand movements but also fine-grained details such as body dynamics, facial expressions, motion trajectories, spatial depth, and body shape, with the goal of fully capturing a person’s pose saunders2020progressive; stoll2020text2sign; pose_format_helper. This end-to-end processing approach greatly simplifies the pose representation Boháček and Hrúz (2022); cho2021unifying; chiang2019unified process that is necessary for SL recognition.
@@
To address these limitations, we develop a recognition framework that operates entirely in the compact pose-rich latent space. Starting from the 2048-dimensional pose features extracted by Vid2Pose, we employ a ResNet34 backbone He et al. (2016) followed by TemporalConv layers Lea et al. (2016) to capture hierarchical temporal patterns. These are jointly learned with the pose features. To refine these predictions into coherent sequences, we apply a Transformer-based encoder-decoder Zhang et al. (2023) that performs sequence-to-sequence refinement with beam search and CTC regularization camgoz2020sign; vaswani2017attention. This design achieves accurate recognition directly in latent space, eliminating the need for raw pose processing and reducing computational overhead compared to vision-based approaches.
Through this muti-stage training, SignX achieves powerful functionality while maintaining architectural simplicity: It learns from various formats of prior knowledge and can directly use the raw video as input. Experimental results show that SignX achieves good performance on our ASL and mainstream datasets Neidle and Metaxas (2025); forster2012rwth; Zhou2021ImprovingSLT-with-monolingual-CSLDaily; WLASL, surpassing existing methods in both accuracy and robustness Chen-arxiv-2023-Robust.
Figure 2: Overview of Constructing Compact Pose-Rich Latent Space: Overall, we utilize ViT Dosovitskiy et al. (2021) to construct and accommodate a pose latent space. It has two entry points: a video entry at the top layer and a pose data entry in the middle section. (a) For training stage 1, we first train the pose fusion layer to output simple text information, ensuring that the learned pose representations are meaningful. (b) For training stage 2, we freeze all other components and only learn how RGB videos can be correctly converted into our pose features. For inference, only RGB is used as input, so we must ensure that we can encode RGB inputs into our pose features.
@@ 2 Related Work
While traditional SLR frameworks predominantly rely on feature extraction from high-dimensional pixel streams or raw skeletal coordinates, these approaches are often hindered by data noise and the difficulty of capturing fine-grained grammatical nuances. In contrast, diffusion models have recently redefined sign language generation by operating within compressed, structured latent spaces Fang et al. (2025b, a). Departing from existing SLR paradigms, our work introduces a novel recognition framework that performs inference directly within a sign-specific latent space. Inspired by the high-efficiency latent modeling in diffusion processes Rombach et al. (2021); Fang et al. (2025d), we leverage a pose-rich latent domain to bypass the limitations of raw data processing. This shift allows our model to more effectively capture complex spatiotemporal dependencies and linguistic structures, establishing a new direction for robust and efficient American Sign Language recognition.
@@ 3.1 Data Processing
To address the challenge of integrating heterogeneous pose information for SL recognition, we first construct a comprehensive processing pipeline that standardizes multiple pose formats. First, we use a high-quality ASLLRP SignStream® 3 Corpus Neidle et al. (2022a); Neidle and Metaxas (2025) dataset, which contains over 80 hours of American Sign Language (ASL) videos with synchronized front view, side view, and facial close-up recordings. For each input video $V\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the number of frames and $H,W$ represent height and width, respectively, we extract sequential pose representations for each pose type: $P_{\text{DWPose}}\in\mathbb{R}^{T\times 384}$, $P_{\text{MediaPipe}}\in\mathbb{R}^{T\times 258}$, $_{\text{SMPLer-X}}\in\mathbb{R}^{T\times 165}, PP$$_{\text{PrimeDepth}}\in\mathbb{R}^{T\times 576}$, $P_{\text{Sapiens}}\in\mathbb{R}^{T\times 576}$. The dimensionality of each pose type reflects its specific information structure. For instance, DWPose contains 18 body keypoints, 21 keypoints for each hand, and 68 facial keypoints, along with their respective confidence scores, totaling 384 dimensions per frame.
@@ 3.2 Step I: ViT for Multimodal Pose Fusion
We initially attempted to use T5 Raffel et al. (2020), currently one of the most popular text decoders, as our text decoding component. However, experimental results revealed that for sign language translation tasks, we only need to consider the vocabulary present in our specific dataset. Large-scale decoders like T5 not only reduce inference speed, but also increase error probability, potentially generating words that do not exist in our target vocabulary $\mathcal{V}_{\text{dataset}}$. Therefore, we developed a custom CodeBook using training data, where the model outputs vocabulary indices rather than raw text tokens, significantly improving accuracy.
Figure 3: Organize and conduct continuous recognition in the latent space: (a) For the latent space of the enriched poses, we further distill, compress and organize it. The number of features trained should be as aligned as possible with the number of Gloss. This can further enhance the upper limit and performance of our model. (b) Then, we develop it based on a BiLSTM method Zhang et al. (2023), enabling it to perform continuous Sign recognition in this latent space, thereby achieving results far superior to previous works.
@@ 3.3 Step II: ViT for Video2Pose
To process raw video inputs directly, we develop a Video2Pose module based on video entry of ViT Dosovitskiy et al. (2021). Given an input video $V\in\mathbb{R}^{T\times H\times W\times 3}$, the Video2Pose module extracts frame-level features and models their temporal relationships:
@@
where $\hat{P}_{i}$ is the predicted pose of type $i$. This end-to-end approach eliminates the need for separate pose extraction models during deployment, greatly simplifying the application process saunders2021mixed; Tarrés et al. (2023); chen2020simple; xie2021segformer; xie2021simmim.
@@ 4.1 Latent Feature Organization and Optimization
Before the temporal learner can reason over sequences, we enforce a disciplined organization of the features produced by the Vid2Pose encoder. Every frame-level embedding $z_{t}\in\mathbb{R}^{2048}$ is accompanied by its gloss supervision signal $g_{t}$ and confidence mask $m_{t}$, giving us a richer tuple $\langle z_{t},g_{t},m_{t}\rangle$. As shown in Fig. 3, this tuple is passed through a lightweight transformer stack that performs (i) variance-preserving layer normalization, (ii) cross-frame covariance whitening (using coefficient $\gamma$), and (iii) stochastic frame dropping (with probability $\rho$) to avoid overfitting to repetitive segments. The process aligns pose cues with textual semantics via a GlossAlign function this process is formalized in Algorithm 1. Here, GlossAlign assigns each feature to its corresponding gloss based on temporal overlap, and CrossCov Zhang et al. (2023) computes the cross-covariance between consecutive frames for redundancy detection.
@@
To combat data scarcity in sign language datasets, we apply a $N=10$-fold temporal augmentation during feature extraction Zhang et al. (2023). For each video sample $i$, the SMKD model Zhang et al. (2023) generates $N$ augmented versions through (i) random temporal scaling $\sigma\in[0.8,1.2]$, (ii) spatial jittering with probability $p_{jit}=0.3$, and (iii) feature-space Gaussian noise $\mathcal{N}(0,0.01)$. The augmented features are stored in HDF5 format with keys $\{i\}_{j}$ where $j\in[0,N-1]$ indexes the augmentation version. It enables efficient random access during training while expanding the effective data size from $M$ to $M\times N$ samples. Unlike pixel-space augmentation which requires re-encoding, our latent augmentation operates on extracted features, reducing computational overhead.
@@ 4.2 Latent-Space Temporal Modeling and Decoding
Sign2Gloss C2SLR Zuo and Mak (2022) 52.50 50.10 38.20 29.50 24.76 52.30 49.95 38.00 29.35 24.58
CorrNet+ Hu et al. (2024) 53.80 51.40 39.50 31.20 26.14 53.60 51.25 39.30 31.05 25.96
SLTUNET Zhang et al. (2023) 54.20 52.10 40.10 31.80 26.48 54.05 51.90 39.95 31.65 26.31
@@
Table 1: Benchmark on the ASLLRP dataset Neidle et al. (2022a); Neidle and Metaxas (2025). To ensure a fair and rigorous comparison on the ASLLRP dataset, we re-implemented and fine-tuned several representative SLR frameworks, including C2SLR and SLTUNET, achieving their best possible performance as modern baselines.
@@ 5 Experimental Results
Here, we evaluate our SignX continuous recognition method on mainstream datasets (ASLLRP Neidle and Metaxas (2025): 2,669 utterances; PHOENIX2014-T camgoz2018neural: 7,096/519/642 train/val/test; CSL-Daily zhou2021improving: 18,401/1,077/1,176; WLASL-2000 WLASL: 2,000 signs), using metrics such as Bleu bleu, WER Klakow and Peters (2002), and P-I Li et al. (2020). We conduct a series of experiments including qualitative, quantitative, ablation, and efficiency evaluations, demonstrating that our method has achieved a significant improvement over the state-of-the-art [^2] More mertric information in Section D.1..
@@ 5.1 Qualitative Assessment
SLTUNET Zhang et al. (2023) 52.23 - - - 27.87 52.11 52.92 41.76 33.99 28.47
TwoStream-SLT Chen et al. (2022) 54.08 54.32 41.99 34.15 28.66 53.48 54.90 42.43 34.46 28.95
CorrNet+ Hu et al. (2024) 54.54 54.56 42.31 34.48 29.13 53.76 55.32 42.74 34.86 29.42
@@
SLTUNET Zhang et al. (2023) 53.58 - - - 23.99 54.08 54.98 41.44 31.84 25.01
TwoStream-SLT Chen et al. (2022) 55.10 55.21 42.31 32.71 25.76 55.72 55.44 42.59 32.87 25.79
CorrNet+ Hu et al. (2024) 55.52 55.64 42.78 33.13 26.14 55.84 55.82 42.96 33.26 26.14
Uni-Sign Li et al. (2025) 56.03 55.30 - - 26.25 56.51 55.08 - - 26.36